I need help to solve this problem, please show work.

If a student attends every management science class, the probability of passing the course is 0.80; but if the student only attends randomly, then the probability of passing the course is 0.50. If a student fails, he or she can take a makeup exam where the probability of passing is 0.60 if the student has attended every class. This probability of passing the makeup exam drops to 0.10 if the student has attended at random. Passing the course is worth 5 credits. Full time attendance ”costs” 3 credits in terms of energy and time, whereas random attendance ”costs” only 1 credit. Use a decision tree to decide which is the best attendance pattern to adopt. Assume that all failing students take the makeup exam and that the payoff for failing is equal to 0.
Answer #1

EMV of daily class = 0.8*5+0.2(0.6*5+0)-3

= 1.6

EMV of random class = 0.5*5+0.5*(0.1*5+0)-1

= 1.75

Random class is better as it has higher EMV

Note: Due to some technical issue, the image could not be uploaded. I will edit the question Later.

0.8 5. a tHends 3 every Class a tteuds raglonely 05 Emu of daily Class= 0.8X5 +o-2 (obxsto)-3 :4706-3 = (.6 EMV OF YandonM C